### v1

# Exploiting Euclidean Distance Field Properties for Fast and Safe 3D planning with a modified Lazy Theta*

Abstract
Graph search planners have been widely used for 3D path planning in the literature, and Euclidean Distance Fields (EDFs) are increasingly being used as a representation of the environment. However, to the best of our knowledge, the integration of EDFs into heuristic planning has been carried out in a loosely coupled fashion, dismissing EDF properties that can be used to accelerate/improve the planning process and enhance the safety margins of the resultant trajectories. This paper presents a fast graph search planner based on a modified Lazy Theta* planning algorithm for aerial robots in challenging 3D environments that exploits the EDF properties. The proposed planner outperforms classic graph search planners in terms of path smoothness and safety. It integrates EDFs as environment representation and directly generates fast and smooth paths avoiding the use of post-processing methods; it also considers the analytical properties of EDFs to obtain an approximation of the EDF cost along the line-of-sight segments and to reduce the number of visibility neighbours, which directly impacts the computation time. Moreover, we demonstrate that the proposed EDF-based cost function satisfies the triangle inequality, which reduces calculations during exploration and, hence, computation time. Many experiments and comparatives are carried out in 3D challenging indoor and outdoor simulation environments to evaluate and validate the proposed planner. The results show an efficient and safe planner in these environments.
keywords:
journal: Robotics and Autonomous Systems
1 Introduction
1.1 Motivation
Path planning in complex 3D environments is a basic procedure that robots must carry out frequently in numerous nowadays applications: tunnel inspection [1], urban firefighting [2], subterranean inspection, like the DARPA Sub-T challenge111 https://www.darpa.mil/program/darpa-subterranean-challenge or drone racing challenges such as AlphaPilot [3]. Approaches ranging from sampling-based to searching-based planners have been successfully applied in the state of the art. We can also find methods based on curve interpolation and optimization. Graph search and heuristic planners such as A* [4] or Theta* [5] are widely used. These methods usually focus on the generation of optimal paths with respect to the path length, leaving aside safety and smoothness of the path (see Fig. 1). As a consequence, computed paths tend to include unnecessary heading changes, which requires post-processing methods [6, 7] to mitigate such effect.
Euclidean Distance Fields (EDF) are a popular environment representation. The field gives the distance to the closest obstacle for any point in space. EDFs have interesting analytical properties [8]. They are particularly useful in 3D path planning [9, 10, 11, 12], but to the best for our knowledge, most existing path planning approaches that integrate EDFs do not exploit their analytical properties. Despite the usefulness of EDF in path planning, heuristic search planners are rarely integrated with them to leverage their properties [13]. Fig. 1 shows the different paths computed by the A* algorithm using a grid map as obstacle representation and the path length as cost, and the A*+M1 [13] which considers an EDF as environment representation and a cost function that integrates costs related to the distance to obstacles. Notice how considering distance to obstacles using EDF as representation enables safer and smoother paths [13]. On the other hand, the path length as cost holds the triangle inequality, which states that the sum of the lengths of two sides of a triangle must be longer than the length of the remaining side. It can be used in heuristic planners to simplify the calculations to select the parent node.
Figure 1: Paths computed by original A* algorithm in blue and red, and computed by A*+M1 from [13] in blue considering distance to obstacles represented as EDF: 2D (left) and 3D (right). Both paths show how the distance to obstacles influences on the computed path by A*+M1.
The goal of this paper is to develop efficient and safe heuristic path planners able to integrate an EDF, consider an EDF-based cost function and take advantage of its properties in order to directly compute safe and smooth paths. It will be also shown that we can even modify existing heuristic planners as both A* and Lazy Theta* planning algorithms to naturally exploit the EDF properties. This will enable approximation of the EDF cost along a segment and reduce the number of visibility neighbours, which affects the computation time. Furthermore, post-processing methods are no longer required to smooth the path. Finally, if we can demonstrate that the proposed EDF-based cost function holds the triangle inequality, the simplification of the calculations will also be ensured. This would be an extension of the previous work [13] because its cost function does not satisfy the triangle inequality.
1.2 Related work
Among the searching-based methods, A* stands as one of the most used solutions. Theta* is a variant of A* that shortens the path length [14], and Lazy-Theta* improves its efficiency [15]. The authors in [5] present a 2D Theta* planner that deal with non-uniform costmaps in 2D, but the paths are not smooth and the triangle inequality is not held, so they have to generalize the original Theta* algorithm and more calculations to select the parent node are performed. Moreover, unnecessary heading changes still occur. In [7] the smoothness of the path is improved with an A* algorithm but they need a post-processing algorithm.
There are also approaches that use traversal cost with Rapidly-exploring Random Trees (RRT)[16] but they also need further post-processing to be smooth. Other approaches compute a path based on heuristic search and linear quadratic minimum-time control but then the initial path has to be refined in a carefully designed B-spline optimization to achieve smooth and safe paths [17]. In other cases, once the initial path is generated, gradient-based methods like CHOMP selects a smoother and safer path [18].
Oleynikova et al. [9] use Euclidean Signed Distance Field (ESDF) to consider the distance cost to the obstacles using Informed RRT*, but the planned trajectory can vary significantly from execution to execution and the EDF properties are not exploited. The authors in [19] adopt a fast marching path planning method on velocity fields using ESDF, but they use Bernstein basis polynomials to generate smooth paths. Moreover, fast marching methods tend to find longer paths than searching-based algorithms as A* since the metric is velocity in the field, not the path length. EDFs are also integrated as environment representation in heuristic planners in the previous work of the authors [13] so that smooth and safe paths are directly generated with a Lazy Theta* algorithm.
1.3 Contributions
The main contributions are:
  • 1.
    A new fast and cost-aware heuristic planner based on a modified Lazy Theta* planning algorithm that integrates an EDF previously computed as environment representation and exploits its properties.
  • 2.
    A new cost function considering the distance to the obstacles so that the triangle inequality is held.
  • 3.
    The computation time of the new EDF-based distance cost term is considerably reduced by leveraging the EDF properties.
  • 4.
    A modified Lazy Theta* algorithm with a novel exploration of the visibility neighbours that foster velocity and reduce the need of post-processing methods.
This paper extends the authors’ previous work [13] by further leveraging the EDF properties, which lead to modifications in the Lazy Theta* algorithm. Additionally, a new cost function is introduced to ensure that the triangle inequality holds. A fully detailed demonstration to approximate the cost along the line-of-sight segments on EDFs is presented, and a new modified Lazy Theta* algorithm is proposed. The exploration of visibility neighbours is significantly reduced by considering the obstacle-gradient information, producing faster, smoother and safer paths than those obtained with [13]. How visibility neighbours are selected also affects safety and smoothness.
The planner is evaluated with many random tests in challenging 3D environments. The implementation of the algorithm and the scripts to launch the tests are open sourced222 https://github.com/robotics-upo/Heuristic_path_planners/.
The paper is organized into five sections. Section II describes the proposed 3D planner. The new cost function and the EDF properties considered are presented in Section III. The selection of the visibility neighbours is described in Section IV. The results and analysis are shown in Section V. Finally, Section VI presents the conclusions.
2 Fast and Safe Path Planner Overview
This section describes the proposed fast and safe path planner (FS-Planner). It is based on the Lazy Theta* algorithm [15], and it leverages the properties of the EDF. Lazy Theta* is a variation of Theta*, which is, in itself, an any-angle algorithm based on A* [20]. Theta* interleaves search and smoothing by including into A* visibility checks that can find new parent nodes if there is a line of sight [14]. In Theta*, the line of sight checks increase the computational load, thus negatively impacting the performance in terms of computation time. The original Lazy Theta* reduces the line of sight checks performed during the search to those strictly necessary. It assumes line of sight between nodes when expanding the search and the triangle inequality is held by the cost function, the path length in this case, delaying the line of sight check. This is relevant in 3D due to the number of visibility neighbours for each node.
The proposed planner integrates an EDF to efficiently represent the robot’s environment and to define the graph employed in the planner. The 3D EDF voxel grid is generated offline considering the distance to the closest occupied node for every node in a map with fixed resolution. As we will see, due to the EDF properties, we can obtain an approximation of the EDF cost along segments (see Section 3) and the EDF representation enables easy obstacle gradient computation based on local EDF derivative (see Section 4).
Algorithm 1 shows the modifications performed in the original Lazy Theta* algorithm (highlighted in red). First of all, a cost function that takes into account the distance to obstacles from the EDF (see $ComputeCost$ function of Algorithm 1) is used with the aim of computing safer paths (the cost function is described in Section 3). Another necessary modification is to limit the line of sight ($LoS$) in the $SetVertex$ function of Algorithm 1 in order to generate smoother paths and foster safety. Otherwise, different nodes of the path could be linked if they are visible to each other, overriding the possible benefits of introducing the distance cost term in order to generate safer paths. These modifications imply an extension of the algorithm rather than a modification on how it works.
1 Main()
2       open := closed := $\emptyset$;
3       g($s_{start}$) := 0;
4       parent($s_{start}$):=$s_{start}$;
5       open.Insert($s_{start}$,$g(s_{start})+h(s_{start})$);
[…]
7             s := open.Pop();
8             SetVertex(s); /* Only Lazy Theta* */
9              If  $s=s_{goal}$ then
10                   return ”path found”;
11
12
13            closed := closed $\cup s$;
14             foreach $s^{\prime}\in nghbr_{vis}(s)$ do
15                   /* Minimum EDF’ */
16                    $s_{edf}$=ComputeMinDerivEDF($s$, $s^{\prime}$);
17
18
19            /* Direction to choose the neighbours */
20              $\hat{u}$=ComputeVector($\hat{s}_{edf}$,$\hat{s}_{goal}$);
21             $nghbr_{edf}(s)$=ChooseNeighbours($\hat{u}$);
22             foreach $s^{\prime}\in nghbr_{edf}(s)$ do
23                   If $s^{\prime}\not\in$ closed then
24                         If $s^{\prime}\not\in$ open then
25                               g($s^{\prime}$) := $\infty$;
26                               parent($s^{\prime}$) := NULL;
27
28
29                        UpdateVertex($s$,$s^{\prime}$);
30
31
32
33
34
35
36      return ”no path found”;
37
38 end
39
40UpdateVertex(s,s’)
41       $g_{old}$ := $g(s^{\prime})$;
42       ComputeCost(s,s’);
43
44      If $g(s^{\prime})<g_{old}$ then
45             If $s^{\prime}\in$ open then
46                   open.Remove(s’);
47
48
49            open.Insert($s^{\prime}$,$g(s^{\prime})+h(s^{\prime})$);
50
51
52
53 end
54ComputeCost(s,s’)
55       /* Path 2 */
56       If $g(parent(s))+||s^{\prime}-parent(s)||+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{c_{w}}{O(parent(s),s^{\prime})}}<g(s^{\prime})$ then
57             $parent(s^{\prime})$ := $parent(s)$;
58             $g(s^{\prime})$ := $g(parent(s))+||s^{\prime}-parent(s)||+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{c_{w}}{O(parent(s),s^{\prime})}}$ ;
59
60
61
62 end
63SetVertex(s)
64       If NOT lineofsight(parent(s),s) OR lineofsight(parent(s),s) $>$ max then
65             /* Path 1 */
66             $parent(s)$ := $\arg\min_{s^{\prime}\in nghbr_{vis}\cap closed}(g(s^{\prime})+||s^{\prime}-s||+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{c_{w}}{O(parent(s),s^{\prime})}})$;
67             $g(s)$ := $\min_{s^{\prime}\in nghbr_{vis}\cap closed}(g(s^{\prime})+||s^{\prime}-s||+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\frac{c_{w}}{O(parent(s),s^{\prime})}})$;
68
69
70
71 end
Algorithm 1 FS-Planner proposed based on modified Lazy Theta*.
The second modification aims to reduce the visibility neighbours to explore around the current node by means of the obstacle-gradient information (see lines 12-17 in Algorithm 1). Depending on how the visibility neighbours are chosen, fast and safe paths can be generated while we explore fewer neighbours. This change does modify the algorithm behavior because the exploration is not performed on all the visibility neighbours as in the original Lazy Theta* algorithm. The selection of the neighbours to explore is explained in Section 4.
3 The cost function
This section presents the new cost function, and how its computation exploits the EDF’s properties. The accumulated cost of a given node in the path in heuristic planners is given by:
$g(s_{i+1})=g(s_{i})+c(s_{i},s_{i+1})$ (1)
where $g(s_{i})$ is the cost of the parent node $s_{i}$ and $c(s_{i},s_{i+1})$ is the cost associated with moving from $s_{i}$ to $s_{i+1}$.
Depending on how you choose $c(s_{i},s_{i+1})$, you can penalize or favour different paths considering EDFs. In our approach, we define it as follows ($ComputeCosts$ in line 34 of Algorithm 1 and $SetVertex$ in line 40 of Algorithm 1):
$c(s_{i},s_{i+1})=||s_{i+1}-s_{i}||+\frac{c_{w}}{O(s_{i},s_{i+1})}$ (2)
where $||s_{i+1}-s_{i}||$ is the Euclidean distance between $s_{i}$ and $s_{i+1}$ (the usual term considered in the original Lazy Theta*), $c_{w}$ is a cost weight factor and $O(s_{i},s_{i+1})$ is the integral of the EDF along the segment defined by $s_{i}$ and $s_{i+1}$.
Notice how the EDF of a node is greater the farther it is from an obstacle; thus, paths are safer when the integral term $O(s_{i},s_{i+1})$ is greater. Considering that the planner minimizes the cost $g$, we inverted the integral of EDF into the cost function as $O(s_{i},s_{i+1})^{-1}$.
Next sections detail the computation of the integral cost over the EDF, and demonstrate that the proposed cost function complies with the triangular inequality.
3.1 Computation of $O(s_{i},s^{\prime}_{i+1})$
The component of the cost containing $O(s_{i},s^{\prime}_{i+1})$ depends on the EDF. The EDF is computed from the distance function to the obstacles, which is defined as:
Definition Let $\Omega\subset\mathbb{R}^{n}$ be a nonempty set. Given any point $x\in\mathbb{R}^{n}$, then the distance from $x$ to the set $\Omega$ is given by the function:
$d(x;\Omega)=inf\{\|x-y\|\;:\;y\in\Omega\}$ (3)
An important property of the distance function is that it is Lipschitz continuous with Lipschitz constant equal to 1 [21, Proposition 1.74], and its gradient norm is $\|\nabla_{x}d(x;\Omega)\|=1$.
The cost $O(s_{i},s^{\prime}_{i+1})$ is defined as the integral of the distance function along the straight segment $l$ that joins $s_{i}$ and $s^{\prime}_{i+1}$ (see Fig. 2):
[…]
where $l(t)$ is the parameterized straight segment between $s_{i}=(x_{i},y_{i},z_{i})$ and $s^{\prime}_{i+1}=(x_{i+1},y_{i+1},z_{i+1})$, defined as:
$l(t)=(x_{i}+(x_{i+1}-x_{i})t,y_{i}+(y_{i+1}-y_{i})t,z_{i}+(z_{i+1}-z_{i})t)^{T},\quad 0\leq t\leq 1$ (5)
so that $l(0)=s_{i}$ and $l(1)=s_{i+1}$. $l^{\prime}(t)$ is the derivative of $l(t)$ with respect to the parameter, which can be easily seen that is $l^{\prime}(t)=s_{i+1}-s_{i}$. Therefore,
[…]
The computation of $O(s_{i},s^{\prime}_{i+1})$ has to be performed repeatedly in Algorithm 1 and it is a computationally expensive step. In the following we derive an approximation to this computation.
Let’s first assume that the obstacles defining the distance field form a convex set. In that case, we have the following:
Proposition The value of $O(s_{i},s^{\prime}_{i+1})$, defined as (4), in the presence of an obstacle represented by a convex set $\Omega$, is bounded as:
$\frac{d(s_{i};\Omega)+d(s_{i+1};\Omega)}{2}||s_{i+1}-s_{i}||-\frac{||s_{i+1}-s_{i}||^{2}}{2}\leq O(s_{i},s^{\prime}_{i+1})\leq\\ \leq\frac{d(s_{i};\Omega)+d(s_{i+1};\Omega)}{2}||s_{i+1}-s_{i}||$ (7)
Proof: First of all, the distance function in $\mathbb{R}^{n}$ with respect to a convex set is a convex function (see [21, Proposition 1.77]). Therefore, $d(l(t);\Omega)$ is a convex function. The Hermite-Hadamard inequality [22] states that if a function $f:[a,b]\rightarrow\mathbb{R}$ is convex the following inequality holds:
[…]
The upper bound of this inequality directly leads to the upper bound in (7).
Regarding the lower bound, we employ the Lipschitz property of the distance function, which states that the distance function at two given points cannot change faster than the distance between the two points. Thus, we have that for the distance function at the mid-point of the segment:
$f(\frac{a+b}{2})=d(l(0.5);\Omega)\geq d(s_{i};\Omega)-\frac{\|s_{i+1}-s_{i}\|}{2}$ (9)
$f(\frac{a+b}{2})=d(l(0.5);\Omega)\geq d(s_{i+1};\Omega)-\frac{\|s_{i+1}-s_{i}\|}{2}$ (10)
and by summing both equations we arrive to the lower bound in (7).
Figure 2: Computation of EDF integration for segments of length $L$ with a convex obstacle. The isocurves of the distance to obstacle are shown in blue.
Given the bounds of (7), we propose to approximate the value of $O(s_{i},s^{\prime}_{i+1})$ by:
$O(s_{i},s^{\prime}_{i+1})\approx\frac{d(s_{i};\Omega)+d(s_{i+1};\Omega)}{2}||s_{i+1}-s_{i}||$ (11)
As mentioned, the bounds for the distance function apply to obstacles that can be represented by a convex set. In general, a scenario will, however, consist of non-convex obstacles. But it is known that a non-convex set can be represented in general as the union of convex sets [23, 24], and each of them will define a Voronoi region in which the distance function $d(l(t);\Omega)$ is convex. Therefore, the approximation in (11) will be valid except for the segments that cross the boundaries between these regions, the cut locus. Therefore, the maximum allowed length of the segments is an important parameter in the method, as longer lengths will lead to potentially larger errors in the approximation, as the probability of crossing the cut locus increases.
Furthermore, note that the difference between the bounds of (7) is $||s_{i+1}-s_{i}||^{2}/2$ or $L^{2}/2$. This demonstration is relevant because the value of $L$ depends on the line of sight considered, that is, the distance between $s_{i}$ and $s_{i+1}$ nodes will always be less than or equal to the maximum line of sight considered. So, the shorter the line of sight or length of the segments, the more similar the bounds are. This will be evaluated in Section 5.
3.2 Triangle inequality
The Lazy Theta* algorithm (and many other heuristic planners) builds on top the triangle inequality [15] that influences on the selection of the parent node. The algorithm selects a parent node of $s_{n}$ between $s_{c}$ and $s_{p}$ (see Figure 3). This selection depends only on whether there is line of sight or not, not the cost, $g$, through them. Therefore, as long as the inequality is satisfied, the path from $s_{p}$ to $s_{n}$ in Fig. 3 can be chosen over path from $s_{p}$ to $s_{n}$ through $s_{c}$ if $s_{n}$ has line of sight to the parent $s_{p}$. If the triangle inequality is not satisfied, both path-costs have to be explicitly computed and compared to select the parent node, which increases the computation time.
Lemma/Proposition Let (2) be the cost function of the FS-Planner. Taking Fig. 3, the triangle inequality is held if the cost from $s_{p}$ to $s_{n}$, $g_{1}(s_{n})$, and the cost $s_{p}$ to $s_{c}$ to $s_{n}$, $g_{2}(s_{n})$, fulfill that:
$g_{1}(s_{n})<g_{2}(s_{n})$ (12)
Proof We can compute $g_{1}(s_{n})$ and $g_{2}(s_{n})$ from (1), (2) and (11) as follows:
$g_{1}(s_{n})=g(s_{p})+L+\frac{2}{R\cdot L}$ (13)
$g_{2}(s_{n})=g(s_{c})+a+\frac{2}{T\cdot a}=g(s_{p})+d+\frac{2}{S\cdot d}+a+\frac{2}{T\cdot a}$ (14)
where $R=P+N^{\prime}$, $S=P+C^{\prime}$, $T=C^{\prime}+N^{\prime}$ and $P,N^{\prime},C^{\prime}$ are the value of the EDF for $s_{p},s_{n},s_{c}$, respectively (see Fig. 3).
Figure 3 shows cases where the obstacle $O1$ is the closest to $s_{p}$, $s_{c}$ and $s_{n}$ or $O2$ is closer than $O1$ to $s_{c}$ and $s_{n}$. We consider the second one because it encompasses all possible cases. The EDF is Lipstchiz continuous, with the Lipstchiz constant equal to $1$, and we can ensure that the difference between the EDF cost of two nodes is always smaller than the distance between the nodes [8]. Therefore, $|P-N^{\prime}|<L$, $|P-C^{\prime}|<d$ and $|N^{\prime}-C^{\prime}|<a$.
Figure 3: Computation of $g(s_{n})$, $s_{n}$ is a neighbour node, through the parent node, $s_{p}$, and the current node, $s_{c}$, considering the distances between nodes and the EDF cost of each node. $L$ is the line of sight or distance between $s_{p}$ and $s_{n}$, $d$ is the distance between $s_{c}$ and $s_{n}$, and $a$ is the resolution of the grid or distance between $s_{c}$ and $s_{n}$. Note that distance can be $a$ or $\sqrt{(}2)\cdot a$
Note that $L<d+a$ due to triangle inequality of the Euclidean distance, so we can discard the terms $g(s_{p})$, $L$, $d$ and $a$ in (13) and (14), and verify that the last term in (13) is smaller than the third and fifth terms in (14) and thus the triangle inequality will be held:
$\frac{2}{R\cdot L}<\frac{2\cdot S\cdot d+2\cdot T\cdot a}{S\cdot T\cdot d\cdot a}$ (15)
This inequality will always be satisfied if the minimum value of the right term is higher than the maximum value of the left term of (15). Taking into account the EDF properties presented above, we obtain $S-T=|P-N^{\prime}|\rightarrow S-T<L$. Therefore, $S$ is upper bounded by $L+T$. We can also consider that $S\cdot T<L\cdot T+T^{2}$ as $S<L+T$. Next we have to check if this value ensures the minimum value of the term on the right of (15). The first derivative of the fraction with respect to $S$ is:
$f^{\prime}(S)=-\frac{2}{S^{2}\cdot d}$ (16)
As the first derivative is negative, the function is decreasing, so the minimum value of (16) is reached with the maximum value of $S$. Regarding the term on the left of (15), we know that $|R-T|<d\rightarrow R=T+d$. In order to consider the smallest value of R, we consider $R=T+a$ as $d\geq a$. Therefore, the inequality follows as:
$\frac{2}{T\cdot L+L\cdot a}<\frac{2\cdot L\cdot d+2\cdot T\cdot d+2\cdot T\cdot a}{L\cdot T\cdot d\cdot a+T^{2}\cdot d\cdot a}$ (17)
In order to calculate the smallest term we have to multiply each numerator by the other denominator:
$2\cdot L\cdot T\cdot d\cdot a+2\cdot T^{2}\cdot d\cdot a<2\cdot L^{2}\cdot T\cdot d+2\cdot T^{2}\cdot L\cdot d\\ +2\cdot L\cdot T^{2}\cdot a+2\cdot L^{2}\cdot d\cdot a+2\cdot L\cdot T\cdot d\cdot a+2\cdot L\cdot T\cdot d\cdot a$ (18)
Note that the term $2\cdot L\cdot T\cdot d\cdot a$ appears in both sides, so it can be simplified. As $L\geq a$, the terms $2\cdot T^{2}\cdot d\cdot a$ and $2\cdot T^{2}\cdot L\cdot d$ can also be simplified. Thus, it is demonstrated that the last term in (13) is always smaller than the third and fifth terms in (14). Therefore, the triangle inequality is held.
4 Selection of visibility neighbours
This section explains how the selection of visibility neighbours from the obstacle-gradient information of the EDF representation and its properties [8] is performed with the aim of generating fast and safe paths. Unlike Lazy Theta*, our approach does not evaluate all possible neighbours around the current node. Instead, we implement a neighbour selection scheme that is intrinsically safe in terms of distance to obstacle, while reducing the number of explored neighbours. For this, we select the neighbours that imply moving away from the obstacle thanks to the obstacle-gradient information and thus safer paths are computed.
First, the directional derivative of the $EDF$, $EDF^{\prime}$, in the direction of each neighbour of the twenty six, $s^{\prime}$, from the current node, $s$, in 3D is computed as follows (see red arrows in Figure 4(a)):
$EDF^{\prime}(s)=\frac{EDF(s)-EDF(s^{\prime})}{distance(s,s^{\prime})}$ (19)
where $EDF(s)$ and $EDF(s^{\prime})$ are the distances to the closest obstacles in each node $s$ and $s^{\prime}$, and $distance$ is the distance between $s$ and $s^{\prime}$. Notice how positive values indicate that we are getting closer to an obstacle in such direction, while negative ones imply moving away from the obstacle.
Then, the minimum value of $EDF^{\prime}(s^{\prime})$ of the twenty six neighbours, $s_{edf}$, is chosen with the $ComputeMinDerivEDF$ function in Algorithm 1 (see red arrow in Figure 4(b)). On the other hand, the direction from each $s^{\prime}$ to the goal is also computed. The direction of the node $s^{\prime}$ with minimum value of $EDF^{\prime}$ to the goal, $s_{goal}$, is the blue arrow in Figure 4(b). The unit vectors of both $s_{edf}$ and $s_{goal}$ give rise to the direction that defines the best candidate node to expand, $\hat{u}$, thanks to the $ComputeVector$ function in line 16 of Algorithm 1 (black arrow is the direction and light gray box the best candidate to expand in Figure 4(b)). Once $\hat{u}$ is computed, the neighbours are selected around the best candidate with the $ChooseNeighbours$ function in line 17 in Algorithm 1 (gray boxes in Figure 4(b)). For example, Figure 4(b) shows nine neighbours selected instead of twenty six neighbours selected in the original algorithm. Finally, the cost, $g$, and parent of each node is computed for the selected neighbours in line 18-23 in Algorithm 1.
This new neighbour selection policy significantly improves search efficiency, as the number of expanded nodes is dramatically decreased. Obviously, the proposed algorithm computes a sub-optimal path because all nodes are not expanded. Nevertheless, experiments will show that the solutions are very close to the optimal. In order to ensure the completeness of the algorithm, if the open list becomes empty [20], the FS-Planner will be run again considering all the neighbours as in [13] to ensure the solution.
Figure 4: Selection of the visibility neighbours considering the obstacle-gradient information and the goal. The black box represents an obstacle.
4.1 Case study
This section presents a study to evaluate the quality of the exploration in the FS-Planner, as it does not guarantee an optimal path because it only expands selected nodes. We will define the quality of selection as the relation between the selected nodes among all the neighbours and the order of these nodes ranked by their estimated cost $g+h$. For simplicity, we analyze the 2D case where eight neighbours are explored. Figure 5 shows two cases by varying the obstacle and goal to consider all possibilities depending on the current node, obstacle, and goal. In the second case, we can analyze cases where the current node is between the obstacle and the goal.
Figure 5: Selection of neighbours in 2D. The current node is in the coordinates’ center, $O_{1}$ and $O_{2}$ are the obstacles, and $G$ the goal. The grid in the centre shows the eight neighbours of the current node. First case (a): $G$ rotates with a small radius around $O_{1}$ (dotted line in red), and $O_{1}$ rotates around the current node (dashed line in red). Second case (b): $G$ rotates around $O_{2}$ (dotted line in blue) with a large radius, and $O_{2}$ rotates around the current node (dashed line in blue). The difference between both setups is that in the second one the current node might fall between $G$ and $O_{2}$.
Let us consider the selection of three and five visibility neighbours out of the eight possible ones in 2D by considering the cases of Fig. 5 and from the obstacle-gradient information (Section 4). In order to measure the quality of the selection, we sort the eight neighbour nodes according to their cost from lowest to highest in each exploration, and analyze which of them are selected. The analysis focuses on determining the percentage of solutions in which at least one of the three lowest-cost nodes is selected among the neighbours selected by our approach: a score of $100\%$ indicates that at least one of the three lowest-cost nodes is in the neighbours selected by our approach in the 100% of the tests. Figure 6 shows the results obtained when three neightbours are selected for the two cases presented in Fig. 5. We can see how the score varies on the line of sight and how, in general, is high, while the minimum score is $79.81\%$. This means that only expanding the three nodes selected by our approach, we are able to get sub-optimal but similar results to expanding all nodes in most cases.
We repeated the same experiment, but selecting five neighbours instead of three. In this case, the score is $100\%$ regardless the line of sight (we do not show in Fig. 6 for simplicity) in both cases. This means that no matter the line of sight, our approach always selects at least one of the lowest-cost nodes if five neighbours are selected. Thus, five neighbours improve the results w.r.t. three, but at the cost of expanding $40\%$ more nodes.
Figure 6: Percent of the quality in the tests performed when three nodes are chosen.
5 Experimental results
This section presents experimental results in different 3D scenarios in order to analyze the robustness and performance of the FS-Planner. The method is benchmarked by considering different number of the visibility neighbours selected among all the nodes. The FS-Planner is tested with different numbers of visibility neighbours in 3D: $9,10,11,13,15$ or $17$ nodes. The proportion of nodes selected relative to all neighbours is similar to the study in 2D ($3$ or $5$ out of $8$), that is, from $9$ to $17$ out of $26$. In all the cases, the neighbours are selected around the best candidate, $\hat{u}$, (see Section 4) except when $10$ nodes are selected. In this particular case, we consider $9$ nodes around the best candidate, $\hat{u}$, and the node opposite to it. A third method is implemented depending on the angle between the minimum gradient unit vector, the unit vectors of $s_{edf}$, and the goal unit vector, the unit vector of $s_{goal}$. If the angle is smaller than $90^{o}$ we consider a number of neighbours and if the angle is greater than $90^{o}$ a larger number of neighbours is considered. For example, $9-11$ considers $9$ neighbours when the angle is smaller than $90^{o}$.
We will benchmark the FS-Planner against the CALT*+M1 planner presented in [13], and the A* algorithms. All planners compute the path on the same EDF previously generated, so the computation times are exclusively referred to the generation of the path. Both FS-Planner and A* make use of the same EDF-based cost function (2). As A*+M1 in [13] mentioned in Section 1, CALT*+M1 considers an EDF as environment representation and a cost function that integrates costs related to the distance to obstacles. The results in [13] showed that the CALT*+M1 provided similar computation times to A*, while smoothness and safety are clearly improved. Moreover, A* algorithm is used as baseline in order to compare them considering the three algorithms the same cost function (2).
The scenarios considered are presented in Fig. 7: S1, like letter H, does not consider segments ”a” and ”c”, and the path goes from 1 to 3 point (see Fig. 7-left); S2, like letter U inverted, it does not consider segments ”b” and ”c” and the path goes from point 2 to 4 (see Fig. 7-left); S3 does not consider segment ”b”, and the path goes from point 2 to 4 (see Fig. 7-left); S4 is the maze shown in Fig. 7-right and the path goes from point 5 to 6; S5 is the same building considered in [13] (see Fig. 9). The objective is to test the FS-Planner in challenging scenarios, with narrow openings and windows considering convex and non-convex obstacles.
Figure 7: Scenarios considered in the tests performed.
The computer used to perform the tests are: Intel Core i7-10700 CPU 2.90GHz, 32 GB RAM, with Ubuntu OS 20.04 LTS. All the planning algorithms have been developed in C++ language and has been integrated into ROS (Robot Operating System) Noetic distribution.
5.1 Dependency on the number of visibility neighbours
This section shows the performance of the FS-Planner depending on the number of visibility neighbours and how it always outperforms CALT*+M1. We focus on the computation time and the explored nodes.
Figure 8 shows the mean computation times and the $95\%$ confidence interval in scenario S1. We consider different values of $c_{w}$ for the CALT*+M1 and the FS-Planner. We can see how the time increases with $cw$ but it remains almost constant from $cw=500$. Notice how the computation time of FS-Planner is always shorter than CALT*+M1 no matter the number of neighbours selected.
Figure 8: Computation time in S1 considering CALT*+M1 and FS-Planner with a different number of selected neighbours: 9, 10, 11, 13, 15 or 17 node. The path goes from 1 to 3 points and $LoS=1$.
Figure 9 shows how the selection of visibility neighbours reduces the explored nodes with the FS-Planner regarding CALT*+M1. The path goes from one building facade to another turning in a corner (the first path considered in scenario S5). Note that the building is not explored inside with the FS Planner when considering the gradient information.
Figure 9: Explored nodes in red to compute a path that goes from one building facade to another turning in a corner with: (left) CALT*+M1 and (right) FS-Planner considering 11 visibility neighbours.
5.2 Random tests
This section presents an evaluation with random tests to compare the CALT*+M1 and the FS-Planner using the A* algorithm as baseline. All planners make use of a previously computed EDF. Both FS-Planner and A* use the same cost function (2). Note that we are interested in the computation times of each planner on EDF and not in the EDF generation. The metrics to evaluate the efficiency, safety and smoothness are as in [13] (see Table 1): path length (length), clearance of obstacles (mean distance related to safety), computation time (time related to efficiency) and heading changes of the path (mean angle related to smoothness). We run twenty path planning tests with different start and goal positions distributed randomly in each scenario. Twenty simulations are performed in each problem to calculate the mean computation time. It is considered $c_{w}=500$ and $LoS=1$ in all the cases. Note that the difference between the bounds of (7) is small ($0.5$), so using the upper bound in (11 provides a good approximation for both convex and non-convex shapes. As a result, we do not need to infer the type of obstacle. The number of visibility neighbours are: 9-11, 10, 11 and 11-13. The solution is always found in the tests shown in this section, but the implementation of the FS-Planner also considers the case that no solution is found. In that case the FS-Planner will be run again considering $17$ neighbours to ensure the solution.
Table 1 summarises the core results in each scenario by using the A* algorithm as baseline, considering (2), in order to compare them. The ratio is calculated as $\frac{algorithm}{A*}$, and the results show the mean of the twenty tests performed and the standard error of the mean which represents the uncertainty about the mean. The best values are highlighted in bold.
All the results show that the time rates of the FS-Planner considering different number of visibility neighbours are lower than the ones of CALT*+M1, as in Fig. 8. Even the FS-Planner shows times lower than A* algorithm over EDF using (2) in scenarios S5 and S6 (see Table 1). Moreover, the explored nodes with CALT*+M1 and the FS-Planner are always lower than the explored nodes with A* algorithm except in S4 that CALT*+M1 is slightly greater (see Table 1-S4). The length is usually very close and similar to the one of A* with CALT*+M1 and FS-Planner. It is greater than the length of A* only with the FS-Planner considering $9-11$ neighbours in S3. This scenario is really challenging and the computed path is longer inside of the U almost closed.
The FS-Planner presents the best values of safety in all scenarios with respect to CALT*+M1 except in S4 in which they are very similar. Also it is safer than A* in scenarios S1, S2, S3 and S6, and very much alike in scenarios S4 and S5. Regarding the smoothness, the FS-Planner and CALT*+M1 compute similar and much smoother paths than A* algorithm, but not in S3 with $9-11$ and $11$ neighbours. Figure 10-left depicts the path computed by A* and the FS-Planner in S4 to show how the smoothness is improved.
It is important to highlight the results obtained in the real scenario considered in [13]. We consider two paths in this real scenario: the first one goes from one building facade to another turning in a corner (see Table 1-S5 and Fig. 10-right); and the second one passes through the window from the outside to the inside (see Table 1-S6). The FS-Planner shows better values than CALT*+M1. Moreover, the computation times of the FS-Planner with 9-11 and 11 neighbours are lower than the ones with A* algorithm. The length and safety are very similar and the paths are much smoother than the ones computed by A* algorithm. Figure 10-right illustrates the improvement of the smoothness with the FS-Planner in S5. Note that the path computed by FS-Planner 9-11 is also smoother than the one computed by A* on the Z axis.
The results demonstrate that the FS-Planner, which exploits the EDF properties and reduces the number of visibility neighbours, works better than CALT*+M1 in terms of computation times, explored nodes, safety and smoothness. Although the length is very similar in both planners, CALT*+M1 computes slightly shorter paths. Regarding A* algorithm over EDF using the cost function (2), the smoothness of the path computed by FS-Planner is considerably improved and the times too in some scenarios. Therefore, we can conclude the FS-Planner improves both CALT*+M1 and A* algorithm over EDF using cost function (2). Moreover, both the FS-Planner and CALT*+M1 do not make use of post-processing methods to achieve smoother paths.
Figure 10: Paths considered in S4 (left) and S5 (right): A* in green and FS-Planner 9-11 in dark green.
Table 1: Evaluation with metrics in the scenarios considered for CALT*+M1 and the FS-Planner considering 9-11, 10, 11, 11-13 and 13 neighbours with $LoS=1$ relative to A*: T is the ratio of the computation time, L is the ratio of the length, N is the ratio of the explored nodes, MD is the ratio of the mean distance to the closer obstacles and MA is the mean angle.
CALT*+M1 FS-Planner9-11 FS-Planner10 FS-Planner11 FS-Planner11-13
S1 T $\downarrow$ $3.653\pm 0.073$ 1.666$\pm$0.053 $2.882\pm 0.046$ $2.297\pm 0.062$ $2.811\pm 0.052$
L $\downarrow$ 0.966$\pm$0.004 $1.020\pm 0.022$ $1.011\pm 0.009$ $1.001\pm 0.013$ $0.990\pm 0.011$
N $\downarrow$ $0.988\pm 0.001$ 0.541$\pm$0.009 $0.9824\pm 0.005$ $0.733\pm 0.014$ $0.867\pm 0.004$
MD $\uparrow$ $0.958\pm 0.017$ $0.884\pm 0.115$ 1.158$\pm$0.119 $0.992\pm 0.071$ $0.943\pm 0.056$
MA $\downarrow$ $0.471\pm 0.030$ $0.532\pm 0.046$ $0.467\pm 0.022$ $0.469\pm 0.049$ 0.443$\pm$0.035
S2 T $\downarrow$ $3.461\pm 0.020$ 1.827$\pm$0.017 $2.830\pm 0.055$ $2.178\pm 0.021$ $2.660\pm 0.025$
L $\downarrow$ 0.974$\pm$0.002 $1.033\pm 0.006$ $0.987\pm 0.006$ $1.019\pm 0.005$ $1.005\pm 0.005$
N $\downarrow$ $0.998\pm 0.001$ 0.634$\pm$0.003 $0.981\pm 0.001$ $0.761\pm 0.002$ $0.877\pm 0.004$
MD $\uparrow$ $1.020\pm 0.038$ 1.050$\pm$0.061 $0.951\pm 0.034$ $0.978\pm 0.049$ $1.008\pm 0.058$
MA $\downarrow$ $0.524\pm 0.014$ $0.576\pm 0.025$ $0.508\pm 0.011$ 0.540$\pm$0.016 $0.459\pm 0.013$
S3 T $\downarrow$ $3.398\pm 0.041$ 1.819$\pm$0.018 $2.660\pm 0.0174$ $2.152\pm 0.016$ $2.619\pm 0.022$
L $\downarrow$ 0.917$\pm$0.001 $1.351\pm 0.012$ $0.980\pm 0.005$ $1.075\pm 0.008$ $1.036\pm 0.007$
N $\downarrow$ $0.998\pm 0.001$ 0.642$\pm$0.004 $0.977\pm 0.002$ $0.779\pm 0.005$ $0.874\pm 0.005$
MD $\uparrow$ $0.971\pm 0.031$ $0.678\pm 0.017$ 1.042$\pm$0.030 $0.796\pm 0.021$ $0.934\pm 0.047$
MA $\downarrow$ 0.530$\pm$0.007 $1.661\pm 0.046$ $0.548\pm 0.021$ $1.099\pm 0.048$ $0.638\pm 0.031$
S4 T $\downarrow$ $3.566\pm 0.020$ 1.328$\pm$0.018 $2.732\pm 0.019$ $2.100\pm 0.017$ $2.547\pm 0.016$
L $\downarrow$ 0.988$\pm$0.004 $0.998\pm 0.006$ $0.989\pm 0.005$ $0.996\pm 0.003$ $0.995\pm 0.002$
N $\downarrow$ $1.004\pm 0.001$ 0.463$\pm$0.004 $0.969\pm 0.002$ $0.738\pm 0.003$ $0.833\pm 0.004$
MD $\uparrow$ 0.991$\pm$0.004 $0.974\pm 0.010$ $0.972\pm 0.010$ $0.969\pm 0.004$ $0.989\pm 0.003$
MA $\downarrow$ $0.588\pm 0.015$ 0.537$\pm$0.007 $0.545\pm 0.018$ $0.566\pm 0.011$ $0.556\pm 0.010$
S5 T $\downarrow$ $2.577\pm 0.033$ 0.782$\pm$0.034 $1.791\pm 0.052$ $0.790\pm 0.034$ $1.699\pm 0.028$
L $\downarrow$ $0.969\pm 0.003$ $0.967\pm 0.009$ $0.962\pm 0.005$ 0.952$\pm$0.004 $0.968\pm 0.004$
N $\downarrow$ $0.997\pm 0.004$ 0.385$\pm$0.014 $0.952\pm 0.005$ $0.406\pm 0.013$ $0.760\pm 0.004$
MD $\uparrow$ $0.989\pm 0.006$ $0.993\pm 0.014$ $0.964\pm 0.005$ $0.973\pm 0.007$ 0.995$\pm$0.007
MA $\downarrow$ $0.430\pm 0.013$ $0.445\pm 0.012$ $0.417\pm 0.009$ $0.414\pm 0.018$ 0.413$\pm$0.016
S6 T $\downarrow$ $2.807\pm 0.029$ 0.789$\pm$0.063 $2.076\pm 0.060$ $0.862\pm 0.119$ $1.722\pm 0.037$
L $\downarrow$ 0.972$\pm$0.043 $1.062\pm 0.027$ $0.996\pm 0.015$ $1.028\pm 0.036$ $0.976\pm 0.011$
N $\downarrow$ $0.999\pm 0.003$ 0.387$\pm$0.033 $0.974\pm 0.026$ $0.427\pm 0.066$ $0.727\pm 0.012$
MD $\uparrow$ $0.987\pm 0.007$ $0.926\pm 0.021$ 1.018$\pm$0.019 $0.948\pm 0.013$ $0.973\pm 0.012$
MA $\downarrow$ $0.670\pm 0.052$ $0.730\pm 0.063$ $0.588\pm 0.037$ $0.624\pm 0.044$ 0.571$\pm$0.033
6 Conclusion and Future Work
A new heuristic search planner based on a modified Lazy Theta* planning algorithm that integrates EDF and exploits its properties has been presented. We proposed the FS-Planner that generates fast and safe paths, and eliminates the need for post-processing methods to smooth the paths. A new cost function has been proposed that adds a cost term related to the distance from obstacles. This provides a safe path planner.
The exploitation of the analytical properties of EDFs have enabled faster computation of the EDF costs along segments. We demonstrate the proposed approximation of $O(s_{i},s^{\prime}_{i+1})$ based on the analytical properties of the EDF (11).Thanks to the EDF properties can also reduce the number of visibility neighbours during path exploration from the obstacle-gradient information. Moreover, the way in which the selection of the visibility neighbors is performed also influences the safety of the path.
Heuristic search planners as Lazy Theta* algorithm take advantage of the triangle inequality to simplify the calculations to select the parent node. We have demonstrated that the EDF-based cost function used in the FS-Planner satisfies the triangle inequality and can be implemented in any heuristic algorithm such as A* or Theta*. Therefore, the FS-Planner improves with respect to CALT*+M1 because it does hold the triangle inequality, resulting in fewer calculations.
These claims have been validated with many tests and comparisons in 3D challenging scenarios and setups. One of the scenarios, S5, is real.
Future work will focus on the integration of new methods for online EDF computation based on neural radiation fields (NeRF). We will analyze how local path planners can take advantage of such online EDFs.
Acknowledgements
This work was partially supported by the grants INSERTION (PID2021-127648OB-C31) and NORDIC (TED2021-132476B-I00), both funded by the “Agencia Estatal de Investigación – Ministerio de Ciencia, Innovación y Universidades” and the “European Union NextGenerationEU/PRTR”.